Find the inverse of f(x) = x/3 + 10
Answer by jim_thompson5910(35256)

f(x) = x/3 + 10

y = x/3 + 10

x = y/3 + 10

x - 10 = y/3

3(x - 10) = y

3x - 30 = y

y = 3x - 30


So the inverse is


If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim